OpenStudy (anonymous):

i really need your help. Which expression can be used to find 73% of 642? 0.73 x 6.42 7.3 x 6.42 0.73 642 7.3 x 642

OpenStudy (anonymous):

73%=73/100=0.73

Parth (parthkohli):

\(73\)% = \(\Large {73 \over 100}\) = 0.73 So, answer is 0.73 * 642

OpenStudy (anonymous):

thankyou so much

OpenStudy (kaederfds):

0,73 \[\times\] 642

Parth (parthkohli):

@cshalvey Wow, lol, again Kaederfds getting the gum :P

OpenStudy (kaederfds):

hi ParthKohli :-)